### v1

# Chromatic numbers, Buchstaber numbers and chordality of Bier spheres

Abstract.
We describe all the Bier spheres of dimension $d$ with chromatic number equal to $d+1$ and prove that all other $d$-dimensional Bier spheres have chromatic number equal to $d+2$, for any integer $d\geq 0$. Then we prove a general formula for complex and mod $p$ Buchstaber numbers of a Bier sphere $\mathrm{Bier}(K)$, for each prime $p\in\mathbb{N}$ in terms of the $f$-vector of the underlying simplicial complex $K$. Finally, we classify all chordal Bier spheres and obtain their canonical realizations as boundaries of stacked polytopes.
Key words and phrases:
Bier sphere, Buchstaber number, chordal graph, chromatic number, stacked polytope
2020 Mathematics Subject Classification:
05C15, 05E45, 13F55, 52B12, 57S12
1. Introduction
This paper is concerned with proving general formulae for the chromatic and Buchstaber numbers of Bier spheres as well as studying their chordality. Our interest in these particular combinatorial invariants and properties is motivated by their relations with each other as well as their applications in topology, algebra and combinatorics.
In the unpublished note [7], any simplicial complex $K$ on $[m]:=\{1,2,\ldots,m\}$ with $m\geq 2$, different from the whole simplex $\Delta_{[m]}$ on $[m]$, was associated with a simplicial complex $\mathrm{Bier}(K)$ defined as the deleted join of $K$ and its Alexander dual $K^{\vee}$. Moreover, it was shown in [7] that $\mathrm{Bier}(K)$ is an $(m-2)$-dimensional PL-sphere with a number of vertices varying between $m$ and $2m$. Since that time this beautiful and simple construction has been studied intensively and found important applications in topological combinatorics [24], polytope theory [19], optimization theory [18], game theory [28], combinatorial commutative algebra [17, 25] and toric topology [21, 22]. New proofs of the fact that it yields a sphere were also obtained, see [23, 24, 25].
Combinatorial and geometrical methods were used to classify all Bier spheres of dimension less than three [21, 22], all flag Bier and Murai spheres [17, 22] and to show that Bier spheres and their generalizations are shellable and edge decomposable [8, 25] as well as that asymptotically the vast majority of them are non-polytopal [24], although no particular examples of a non-polytopal Bier sphere have been found so far. Furthermore, methods of homological algebra allowed us to compute certain multigraded Betti numbers of Bier spheres and their generalizations [17, 25]. Finally, the application of topological combinatorics methods showed that all Bier spheres are starshaped [18], and methods of toric topology yielded the classification of all minimally non-Golod Bier spheres [22].
The structure of the paper is as follows. In Section 2, we recall the key definitions and introduce basic examples related to the notions of a simplicial complex, simple graph, convex polytope, and Bier sphere.
The first combinatorial invariant of a Bier sphere studied in this paper is the chromatic number discussed in Section 3. By definition, this is the least number of colors needed for a proper coloring of the vertex set of a simplicial complex; that is, vertices of each simplex are colored in pairwise different colors. The chromatic number is one of the most important classical combinatorial invariants of a simple graph as well as of a simple polytope. By definition, the chromatic number $\chi(P)$ of an $n$-dimensional convex simple polytope $P$ equals that of its nerve complex $\partial P^{*}$ and hence $m\geq\chi(P)\geq n$, where $m$ is the number of facets of $P$. A combinatorial description of $n$-colorable simple $n$-polytopes was obtained in [20], while that in the case of $(n+1)$-colorable simple $n$-polytopes remains unknown.
Recall that a triangulated sphere is called polytopal if it is simplicially isomorphic to the boundary of a convex simplicial polytope. We prove that for any simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$ the chromatic number $\chi(\mathrm{Bier}(K))$ equals either $m-1$, or $m$. More precisely, the following statement holds, see Theorem 3.16.
Theorem 1.1.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then
  1. (a)
    $m-1\leq\chi(\mathrm{Bier}(K))\leq m$;
  2. (b)
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next two cases holds:
    • $P_{K}=I^{m-1}$;
    • $P_{K}=I^{m-3}\times P_{6}$.
Here, we denote by $P_{K}$ a simple polytope such that for a polytopal Bier sphere $\mathrm{Bier}(K)$ one has a combinatorial equivalence $\mathrm{Bier}(K)=\partial P_{K}^{*}$. We also use the notations $I^{n},n\geq 0$ for the standard $n$-dimensional cube and $P_{m},m\geq 3$ for an $m$-gon. Moreover, Lemma 3.15 shows that for $m\geq 3$, the chromatic number $\chi(\mathrm{Bier}(K))=m-1$ if and only if $K$ is obtained starting from one of the following three simplicial complexes on the set $[3]$: either (i) two points forming a segment complex, or (ii) three points forming a chain complex, or (iii) three disjoint points complex by iterative application of taking a cone and taking Alexander dual operations.
The second combinatorial invariant of a Bier sphere studied in this paper is the Buchstaber number discussed in Section 4. By definition, the (complex) Buchstaber number of a simplicial complex $K$ with $m$ vertices is the maximal integer $r$ such that there exists a characteristic map $\Lambda\colon[m]\to\mathbb{Z}^{m-r}$; that is, each simplex of $K$ is mapped to a part of a lattice basis of $\mathbb{Z}^{m-r}$. We denote it by $s(K)$. Replacing $\mathbb{Z}$ by $\mathbb{Z}_{p}$ for a prime $p\in\mathbb{N}$ and the integer lattice by a $\mathbb{Z}_{p}$-vector space above, we get a definition of the mod $p$ Buchstaber number of $K$. We denote it by $s_{p}(K)$. The theory of complex and real Buchstaber numbers has been developed in the framework of toric topology and combinatorial commutative algebra, see [1, 2, 12, 13, 4, 5].
These invariants play a crucial role in toric topology [9, 10] as they reflect important topological properties of polyhedral products [3]. Namely, the complex Buchstaber number of $K$ equals the maximal rank of a toric subgroup $H$ in the compact torus $(\mathbb{S}^{1})^{m}\subset({\mathbb{D}}^{2})^{m}$ such that the restricted coordinatewise action of $H$ on the moment-angle-complex $\mathcal{Z}_{K}\subseteq({\mathbb{D}}^{2})^{m}$ is free. Similarly one can interpret the mod $p$ Buchstaber numbers, see [6].
It turns out that chromatic and Buchstaber numbers of an $(n-1)$-dimensional simplicial complex $K$ with $m$ vertices are closely related to each other by the following inequalities:
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n,$
for any prime $p\in\mathbb{N}$, see [6, Corollary 5.8]. The real ($p=2$) and complex Buchstaber invariants of $\mathrm{Bier}(K)$ were computed in [21, Theorem 3.4] in the assumption that all the ghost vertices of $K$ are taken into account. In  Theorem 4.2 we get rid of that restriction proving the following result, which provides a general formula for all the Buchstaber numbers of $\mathrm{Bier}(K)$ only in terms of the $f$-vector of $K$.
Theorem 1.2.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then we have
$s(\mathrm{Bier}(K))=s_{p}(\mathrm{Bier}(K))=f_{0}(K)-f_{m-2}(K)+1,\text{ for any prime }p\in\mathbb{N}.$
Since $\dim(\mathrm{Bier}(K))=m-2$ for any simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$, polytopal Bier spheres give rise to $(m-1)$- and $m$-colorable $(m-1)$-dimensional simple polytopes, which admit characteristic maps and therefore quasitoric manifolds and small covers due to [11].
Finally, chordality of Bier spheres is discussed in Section 4. Recall that a simple graph is chordal if it contains no induced subgraphs being cycles of length greater than three. We call a simplicial complex chordal if its 1-skeleton is a chordal graph. It turns out that this combinatorial property acquires algebraic and topological interpretations in the framework of toric topology.
[…]
In Theorem 4.7 we classify all chordal Bier spheres by proving the following result.
Theorem 1.3.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then $\mathrm{Bier}(K)$ is chordal if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next three cases holds:
  • $P_{K}=\Delta^{1}$, where $m=2$;
  • $P_{K}=\Delta^{2}$, where $m=3$;
  • $P_{K}=\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
Here, by $\mathrm{vc}^{k}(\Delta^{m-1})$ we mean a truncation polytope; that is, a simple polytope obtained from $\Delta^{m-1}$ by a sequence of $k$ vertex cuts by hyperplanes in general position. Moreover, as Lemma 4.6 asserts, in case $m\geq 4$, chordality of $\mathrm{Bier}(K)$ is equivalent to either $K$ or $K^{\vee}$ having no edges. This yields a canonical realization for the corresponding Bier sphere as a boundary of a stacked polytope; that is, one gets $\mathrm{Bier}(K)$ in that case as a boundary of a simplicial $(m-1)$-dimensional polytope obtained from the simplex $\Delta^{m-1}$ by performing stellar subdivisions (glueing pyramids) over some of its facets.
2. Basic definitions and constructions
In this section, we introduce our notations and recall some necessary definitions, examples and results related to the Bier sphere construction as well as to the notion of a chromatic number for simple graphs and, more generally, for simplicial complexes.
Definition 2.1.
Set $[m]:=\{1,2,\ldots,m\}$ for an integer $m\geq 1$. We say that a non-empty subset $K\subseteq 2^{[m]}$ is an abstract simplicial complex on $[m]$ if the following condition holds:
$\sigma\in K,\tau\subseteq\sigma\Rightarrow\tau\in K.$
Elements of $K$ are called its faces or simplices. Faces of cardinality $1$ are called (geometrical) vertices, and faces of cardinality $2$ are called edges.
The set of vertices of $K$ will be denoted by $V(K)$. Maximal (by inclusion) simplices of $K$ are called its facets, and their set is denoted by $\mathrm{M}(K)$; we say that facets of $K$ generate $K$.
Minimal (by inclusion) subsets in $[m]$ that are not simplices of $K$ are called its minimal non-faces, and their set is denoted by $\mathrm{MF}(K)$. Minimal non-faces of cardinality $1$ are called ghost vertices.
Recall that the dimension of a simplex is one less than its cardinality; the dimension of a simplicial complex $K$ equals the maximal dimension of its simplex and is denoted by $\dim(K)$. If all maximal faces of $K$ have the same cardinality, then $K$ is called pure.
One of the most important combinatorial invariants (i.e. characteristics invariant under a face lattice isomorphism) of a simplicial complex is its face vector. Namely, the $f$-vector of a simplicial complex $K$ on $[m]$ is a tuple $f(K):=(f_{-1}(K),\ldots,f_{m-1}(K))$, where $f_{i}(K)$ denotes the number of faces of dimension $i$ in $K$. Note that $f_{-1}(K)=1$ for any simplicial complex $K$, since $\varnothing\in K$ is the only simplex in $K$ of dimension $-1$, and $f_{i}(K)=0$ if $i>\dim(K)$.
Example 2.2.
Let $P$ be a convex simple $n$-dimensional polytope with $m$ facets; that is, a bounded intersection of $m$ closed halfspaces in $\mathbb{R}^{n}$ being in general position. Then the boundary $\partial P^{*}$ of its dual simplicial polytope $P^{*}$ is a geometric realization of a certain abstract simplicial complex which we denote by $K_{P}$ and call the nerve complex of $P$. Thus, we have:
$f_{0}(K_{P})=m\text{ and }\dim(K_{P})=n-1.$
Example 2.3.
Let $m\geq 3$ and $P_{m}\subset{\mathbb{R}}^{2}$ be an $m$-gon. Then its dual polytope $P_{m}^{*}$ is isomorphic to $P_{m}$. Hence, $P_{m}$ is a simple and a simplicial polytope at the same time. The nerve complex of $P_{m}$ equals its boundary $\partial P_{m}$, which is the $m$-cycle denoted in what follows by $Z_{m}$.
$Z_{3}$
$Z_{4}$
$Z_{5}$
$Z_{6}$
Figure 1. The only cycles that are 1-dimensional Bier spheres.
Example 2.4.
Consider the following simplicial complexes on $[m]$ with $m\geq 1$:
  • $\varnothing_{[m]}:=\{\varnothing\}$, the void complex;
  • $\Delta_{[m]}:=2^{[m]}$, the simplex;
  • $\partial\Delta_{[m]}:=2^{[m]}\setminus\{[m]\}$, the boundary of the simplex.
Note that $\dim(\varnothing_{[m]})=-1$, $\dim(\Delta_{[m]})=m-1$, and $\dim(\partial\Delta_{[m]})=m-2$, for any $m\geq 1$. Moreover, all the three simplicial complexes are pure and $\mathrm{MF}(\Delta_{[m]})=\varnothing$, $\mathrm{MF}(\partial\Delta_{[m]})=\{[m]\}$, while $\varnothing_{[m]}$ has $m$ ghost vertices.
In what follows, we sometimes interchangeably use the notations $i$ and $\{i\}$ for a (ghost) vertex of a simplicial complex $K$ on $[m]$ when $i\in[m]$.
Definition 2.5.
Let $K$ be a simplicial complex on $[m]$ and $I\subseteq[m]$. We call the simplicial complex
$K_{I}:=\{J\in K\,|\,J\subseteq I\}=K\cap 2^{I}$
the full subcomplex in $K$ on the set $I$. In particular, if $i\in[m]$, then the deletion of $i$ from $K$ is the full subcomplex $\operatorname{Del}_{K}(i)$ in $K$ on the set $[m]\setminus\{i\}$. If $I\in K$ we call the simplicial complex
$\operatorname{Link}_{K}(I):=\{J\in K\,|\,I\cup J\in K,I\cap J=\varnothing\}$
the link in $K$ of the face $I$.
Example 2.6.
Let $K$ be a simplicial complex on $[m]$. Observe that the following equivalences hold:
  • $I\in K$ if and only if $K_{I}=\Delta_{I}$;
  • $I\in\mathrm{MF}(K)$ if and only if $K_{I}=\partial\Delta_{I}$;
  • $\operatorname{Link}_{K}(I)=K$ if and only if $I=\varnothing$;
  • $\operatorname{Link}_{K}(I)=\{\varnothing\}$ if and only if $I\in M(K)$.
A particularly important class of simplicial complexes is formed by all (simple) graphs. Throughout this paper, by a graph $\Gamma=\Gamma(V,E)$ with the vertex set $V$ and the edge set $E\subseteq\binom{V}{2}$ we mean a simplicial complex on $V$ of dimension not greater than $1$ and without ghost vertices.
The $n$-skeleton $\mathrm{sk}^{n}(K),n\geq 0$ of a simplicial complex $K$ is a simplicial complex with no ghost vertices consisting of all simplices of $K$ of dimension not greater than $n$. Note that both $\mathrm{sk}^{0}(K)$ and $\mathrm{sk}^{1}(K)$ are graphs. Moreover, one has:
$V(\mathrm{sk}^{0}(K))=V(K)=V(\mathrm{sk}^{1}(K))\text{ and }\mathrm{sk}^{0}(K)=V(K)\sqcup\{\varnothing\},$
for any simplicial complex $K$.
Here is a definition of a simplicial operation, which will be used frequently in what follows.
Definition 2.7.
Let $K_{1}$ be a complex on $V_{1}$ and $K_{2}$ be a complex on $V_{2}$ with $V_{1}\cap V_{2}=\varnothing$. Then we define their join $K$ as a simplicial complex on $V_{1}\sqcup V_{2}$ such that
$K=K_{1}\ast K_{2}:=\{\sigma_{1}\sqcup\sigma_{2}\,|\,\sigma_{i}\in K_{i},\,i=1,2\}.$
In particular, we have
  • $\operatorname{Cone}_{v}(K):=\{v\}\ast K$, the cone over $K$ with apex $v$;
  • $\Sigma_{v,w}(K):=\partial\Delta_{\{v,w\}}\ast K$, the suspension over $K$ with vertex pair $\{v,w\}$.
In particular, if $K=\operatorname{Cone}_{v}(L)$ is a cone over $L$ with apex $v$, then $L=\operatorname{Del}_{K}(v)$ is a deletion of $v$ from $K$.
Definition 2.8.
Let $\Gamma=\Gamma(V,E)$ be a graph with the vertex set $V$ and the edge set $E$. A proper coloring of $\Gamma$ is a surjective map $c\colon V\to C$ such that
$\text{if}\quad\{i,j\}\in E,\quad\text{then}\quad c(i)\neq c(j).$
The chromatic number $\chi(\Gamma)$ of the graph $\Gamma$ is the smallest cardinality of the set of colors $C$ such that there exists a proper coloring $c\colon V\to C$. We set the chromatic number $\chi(K)$ of a simplicial complex $K$ to be equal to $\chi(\mathrm{sk}^{1}(K))$ and we say that $c$ is a proper coloring of $K$ if it is a proper coloring of $\mathrm{sk}^{1}(K)$. Finally, we set the chromatic number $\chi(P)$ of a simple polytope $P$ to be equal to $\chi(K_{P})$ and, similarly, we introduce the notion of a proper coloring of $P$.
The next result is a direct corollary from the definition of a chromatic number and is well-known.
Proposition 2.9.
The following statements hold.
  • (a)
    Let $m\geq 3$. Then one has:
    $\chi(P_{m})=2,\text{ if $m$ is even and }\chi(P_{m})=3,\text{ if $m$ is odd.}$.
  • (b)
    For any simplicial complex $K$ and a disjoint point $v$ one has:
    $\chi(\operatorname{Cone}_{v}(K))=\chi(K)+1.$
  • (c)
    For any simplicial complex $K$ and two distinct points, $v$ and $w$, disjoint from $K$ one has:
    $\chi(\Sigma_{\{v,w\}}(K))=\chi(K)+1.$
Now we are going to introduce the main object of study in this paper, a Bier sphere. In the definition below, we use the following notations. Let $[m]:=\{1,2,\ldots,m\}$ and $[m^{\prime}]:=\{1^{\prime},2^{\prime},\ldots,m^{\prime}\}$ be two disjoint ordered sets with the map $\phi\colon i\mapsto i^{\prime},1\leq i\leq m$ being an order preserving bijection between them. Denote by $I^{\prime}$ the image $\phi(I)$ of a subset $I\subseteq[m]$.
Definition 2.10.
Suppose $K\neq\Delta_{[m]}$ is a simplicial complex on $[m]$ with $m\geq 2$. Then
  • the Alexander dual of $K$ is a simplicial complex $K^{\vee}$ on $[m^{\prime}]$ such that
    $I\in\mathrm{MF}(K)\Longleftrightarrow[m^{\prime}]\setminus I^{\prime}\in\mathrm{M}(K^{\vee});$
  • the Bier sphere of $K$ is a simplicial complex $\mathrm{Bier}(K)$ on $[m]\sqcup[m^{\prime}]$ such that
    $\mathrm{Bier}(K)=\{I\sqcup J^{\prime}\,|\,I\in K,J^{\prime}\in K^{\vee},I\cap J=\varnothing\};$
    that is, $\mathrm{Bier}(K)$ is the deleted join of $K$ and $K^{\vee}$.
It was shown in [7] that $\mathrm{Bier}(K)$ defined above is a PL-sphere of dimension $(m-2)$ with the number of vertices varying between $m$ and $2m$. It follows from the definition that the Bier sphere construction is symmetric: $\mathrm{Bier}(K)=\mathrm{Bier}(K^{\vee})$. In Figure 1 one can find the geometric realizations of all possible 1-dimensional Bier spheres, see also [22, Example 2.8].
Example 2.11.
Let $K_{1}=\Delta_{[m-1]}$ and $K_{2}=\Delta_{[m-2]}$ be simplicial complexes on $[m]$, where $m\geq 3$. Then $K_{1}^{\vee}=\Delta_{[(m-1)^{\prime}]}$ and $K_{2}^{\vee}=\Delta_{[(m-1)^{\prime}]}\cup\Delta_{[(m-2)^{\prime}]\cup\{m^{\prime}\}}$. Observe that a maximal simplex in $\mathrm{Bier}(K_{1})$ is of a form $([m-1]\setminus I)\sqcup I^{\prime}$, where $I\subseteq[m-1]$, hence
$\mathrm{Bier}(K_{1})=\partial\Delta_{\{1,1^{\prime}\}}\ast\partial\Delta_{\{2,2^{\prime}\}}\ast\cdots\ast\partial\Delta_{\{m-2,(m-2)^{\prime}\}}\ast\partial\Delta_{\{m-1,(m-1)^{\prime}\}}.$
Furthermore, a maximal simplex in $\mathrm{Bier}(K_{2})$ is of a form $([m-2]\setminus I)\sqcup I^{\prime}\sqcup\{k^{\prime}\}$, where $I\subseteq[m-2]$ and $k\in\{m-1,m\}$, hence
$\mathrm{Bier}(K_{2})=\partial\Delta_{\{1,1^{\prime}\}}\ast\partial\Delta_{\{2,2^{\prime}\}}\ast\cdots\ast\partial\Delta_{\{m-2,(m-2)^{\prime}\}}\ast\partial\Delta_{\{(m-1)^{\prime},m^{\prime}\}}.$
Thus, we see that both $\mathrm{Bier}(K_{1})$ and $\mathrm{Bier}(K_{2})$ are $(m-3)$-times iterated suspensions over the 4-cycle $Z_{4}$, or, in other words, both are isomorphic to the boundary of the $(m-1)$-dimensional cross-polytope $(I^{m-1})^{*}$.
Taking into account Proposition 2.9 and Example 2.11, our first goal is to study the chromatic numbers of Bier spheres. When $m=2$, the only Bier sphere has a geometric realization as the set of two disjoint points, it has chromatic number $1$. Therefore, in what follows, if otherwise not stated explicitly, we always assume that a simplicial complex on $[m]$ under consideration is different from $\Delta_{[m]}$ and that $m$ is always greater or equal to $3$.
When $m=3$, it is clear that the Bier sphere $\mathrm{Bier}(K)$ is a boundary of a polygon, and so its chromatic number is equal to either $2$ or $3$, depending on the parity of the number of vertices, see Proposition 2.9 (a). Up to isomorphism and taking Alexander dual, we have only five simplicial complexes on [3] shown in Figure 2.
$\Gamma_{3}$
$\Gamma_{4}$
$G_{4}$
$\Gamma_{5}$
$\Gamma_{6}$
Figure 2. The complexes $G_{4}$ and $\Gamma_{6}$ are self-dual and $G_{4}$ has a ghost vertex. The indices in the notation of a complex indicate the number of vertices in the corresponding Bier sphere.
The chromatic number of Bier spheres $\mathrm{Bier}(\Gamma_{4})$, $\mathrm{Bier}(G_{4})$, and $\mathrm{Bier}(\Gamma_{6})$ is 2, since it is $Z_{4}$ in the first two cases and $Z_{6}$ in the last case. The chromatic number of $\mathrm{Bier}(\Gamma_{3})$ and $\mathrm{Bier}(\Gamma_{5})$ is 3, since they are isomorphic to $Z_{3}$ and $Z_{5}$, respectively. This motivates us to compute the chromatic number of an arbitrary Bier sphere: it is done in the next section.
3. Bier spheres and their chromatic numbers
In this section, we show that every Bier sphere $\mathrm{Bier}(K)$ of a simplicial complex $K$ on $[m]$ with $m\geq 2$ has the chromatic number equal to either $m-1$, or $m$ and describe all the simplicial complexes $K$, for which $\chi(\mathrm{Bier}(K))=m-1$. The discussion of chromatic numbers of 1-dimensional Bier spheres in the end of the previous section as well as Example 2.11 will serve as the first step in the desired description.
Proposition 3.1.
Let $K$ be a simplicial complex on $[m]$. Then
$\max(m-1,\chi(K),\chi(K^{\vee}))\leq\chi(\mathrm{Bier}(K))\leq m.$
Proof.
Since $\dim(\mathrm{Bier}(K))=m-2$, the chromatic number $\chi(\mathrm{Bier}(K))\geq m-1$. On the other hand, the map
$i,i^{\prime}\mapsto c_{i},\text{ for }1\leq i\leq m,$
yields a proper coloring of the vertex set of $\mathrm{Bier}(K)$ by the set of colors $C=\{c_{1},\ldots,c_{m}\}$, hence $\chi(\mathrm{Bier}(K))\leq m$.
Finally, since both $K$ and $K^{\vee}$ are full subcomplexes in $\mathrm{Bier}(K)$ by definition of a Bier sphere, one has
$\chi(\mathrm{Bier}(K))\geq\chi(K),\chi(K^{\vee}),$
which finishes the proof. ∎
Next, we are going to introduce two classes of simplicial complexes that contain the classes of cones and suspensions, see Definition 2.7, respectively.
Definition 3.2.
We say that a graph $\Gamma$ with the vertex set $V$ is
  • a cone graph if there exists $a\in V$ such that:
    $\operatorname{Link}_{\Gamma}(a)=V\setminus\{a\}.$
    We call a simplicial complex $K$ a weak cone if its 1-skeleton $\mathrm{sk}^{1}(K)$ is a cone graph;
  • a suspension graph if there exist distinct $a,b\in V$ such that:
    $\operatorname{Link}_{\Gamma}(a)=V\setminus\{a,b\}=\operatorname{Link}_{\Gamma}(b).$
    We call a simplicial complex $K$ a weak suspension if its 1-skeleton $\mathrm{sk}^{1}(K)$ is a suspension graph.
Obviously, any cone over a simplicial complex is a weak cone complex, and any suspension over a simplicial complex is a weak suspension complex. The opposite implications fail, as the following example illustrates.
Example 3.3.
Any geometric realization of a simplicial sphere is a topological suspension. On the other hand, the boundary of an octahedron is a suspension and, in particular, a weak suspension with respect to each pair of its opposite vertices, whereas both the boundaries of a tetrahedron and an icosahedron are not even weak suspensions, see Figure 3.
Figure 3. Octahedron is a suspension and therefore a weak suspension, but tetrahedron and icosahedron are not even weak suspensions.
Here is a family of simplicial spheres, one sphere for each $m\geq 5$, such that each member of the family is a weak suspension, but not a suspension.
Example 3.4.
Let $m\geq 5$ and consider the following simplicial complex $K_{m}$ on the vertex set $[m]$:
$\mathrm{MF}(K_{m})=\{\{1,2,\ldots,m-2\},\{2,3,\ldots,m-1\},\{2,m\},\ldots,\{m-1,m\}\}.$
By definition, its Alexander dual complex $K_{m}^{\vee}$ is a simplicial complex on the vertex set $[m^{\prime}]$ given by
[…]
Thus, $K_{m}^{\vee}$ is a union of two edges and the cone with apex $1^{\prime}$ over the standard triangulation of $\partial\Delta_{\{2^{\prime},\ldots,(m-1)^{\prime}\}}$. This implies that the Bier sphere $\mathrm{Bier}(K_{m})$ is a weak suspension with respect to the unique vertex pair $\{1,1^{\prime}\}$. On the other hand, neither $K_{m}$ nor $K_{m}^{\vee}$ is a cone, so $\mathrm{Bier}(K_{m})$ is not a suspension.
12345
$2^{\prime}$$4^{\prime}$$3^{\prime}$$1^{\prime}$$5^{\prime}$
123456
$1^{\prime}$$2^{\prime}$$3^{\prime}$$4^{\prime}$$5^{\prime}$$6^{\prime}$
Figure 4. On the left it is the simplicial complex $K_{5}$ and its Alexander dual $K_{5}^{\vee}$. On the right it is the simplicial complex $K_{6}$ (there is no face $\{2,3,4,5\}$) and its Alexander dual $K_{6}^{\vee}$.
Proposition 3.5.
Let $m\geq 5$. Then both $K_{m}$ and $K_{m}^{\vee}$ are weak cones, and one has:
$\chi(K_{m})=\chi(K_{m}^{\vee})=m-1.$
On the other hand, $\chi(\mathrm{Bier}(K_{m}))=m$.
Proof.
Consider the following coloring $c\colon[m]\sqcup[m^{\prime}]\to C(m)=\{c_{1},\ldots,c_{m-1}\}$ defined by
$c(i)=c(i^{\prime})=c_{i}\text{ for }1\leq i\leq m-1\text{ and }c(m)=c(m^{\prime})=c_{2}.$
It is easy to see that $C(m)$ provides a proper coloring of both $K_{m}$ and $K_{m}^{\vee}$ with $(m-1)$ colors, and so
$\chi(K_{m}),\chi(K_{m}^{\vee})\leq m-1.$
On the other hand, the 1-skeleta of the full subcomplexes $K_{m}\cap 2^{[m-1]}$ and $K^{\vee}_{m}\cap 2^{[(m-1)^{\prime}]}$ are complete graphs with $(m-1)$ vertices, and therefore
$\chi(K_{m}),\chi(K_{m}^{\vee})\geq m-1.$
Combining the last two inequalities above yields the first statement. The second statement follows from Proposition 3.1 and the observation that the 1-skeleta of the full subcomplexes
$\mathrm{Bier}(K_{m})\cap 2^{[m-1]\cup\{m^{\prime}\}}\text{ and }\mathrm{Bier}(K_{m})\cap 2^{[(m-1)^{\prime}]\cup\{m\}}$
are complete graphs with $m$ vertices. ∎
Skeleta of a simplex never yield a Bier sphere, which is a suspension, as shown in the following example. In the case of the 0-skeleton of the $m$-simplex and its Alexander dual, which is the $(m-3)$-skeleton of the $m$-simplex, the Bier sphere is not even a weak suspension, since it does not contain any ghost vertices, nor does it contain both edges $\{i,j\}$ and $\{i^{\prime},j^{\prime}\}$, for each pair of distinct vertices $i,j\in[m]$. The next example shows that in all other cases, the Bier sphere of a skeleton of a simplex is a weak suspension and not a suspension.
Example 3.6.
Let $m\geq 5$, $41\leq k\leq m-$ 4, and $K=\Delta_{[m]}^{(k)}$, i.e., the set of maximal simplices is $\mathrm{M}(K)=\{I\subset[m]\,|\,|I|=k+1\}$. Then $I^{\prime}$ is a minimal nonsimplex in $K^{\vee}$ if and only if $|I|=m-k-2$, hence $K^{\vee}=\Delta_{[m^{\prime}]}^{(m-k-3)}$. For every $i,j\in[m]$, $i\neq j$, we have $\{i,j^{\prime}\}$, $\{i^{\prime},j\}\in\mathrm{Bier}(K)$, since there are no ghost vertices in $K$ and $K^{\vee}$, $\{i,j\}\in K\subset\mathrm{Bier}(K)$, since $k\geq 1$, and $\{i^{\prime},j^{\prime}\}\in K^{\vee}\subset\mathrm{Bier}(K)$, since $m-k-3\geq 1$. Therefore $\mathrm{Bier}(K)$ is a weak suspension for every pair $\{i,i^{\prime}\}$. For every $i\in[m]$ we can write $[m]\setminus\{i\}=I\sqcup J$, where $|I|=i+1$ and $|J|=m-i-2$, hence $I\sqcup J^{\prime}$ is a maximal simplex in $\mathrm{Bier}(K)$, so $\mathrm{Bier}(K)$ is not a suspension with a vertex pair $\{i,i^{\prime}\}$ and therefore it is not a suspension. Finally, all the three chromatic numbers $\chi(K)$, $\chi(K^{\vee})$, and $\chi(\mathrm{Bier}(K))$ are equal to $m$, since the complete graph on $m$ vertices is a subcomplex of all the three complexes.
The next result is formally a generalization of Proposition 2.9 (b) and (c). However, in fact two statements are equivalent, since chromatic number of a simplicial complex depends only on its 1-skeleton.
Lemma 3.7.
Suppose a simplicial complex $L$ is a weak suspension or a weak cone over a simplicial complex $K$. Then we have:
$\chi(L)=\chi(K)+1.$
Proof.
Note that one can add one color to any coloring of $K$ and color an apex (vertex pair) of the weak cone (weak suspension) $L$ over $K$ in that new color. This shows that
$\chi(L)\leq\chi(K)+1.$
On the other hand, for any coloring $C$ of $L$, the apex (vertex pair) of the weak cone (weak suspension) $L$ over $K$ is colored in colors not belonging to the restriction $C|_{K}$. This implies that
$\chi(K)<\chi(L).$
Combining the two inequalities above yields $\chi(L)=\chi(K)+1$, which finishes the proof. ∎
Lemma 3.8.
Let $K$ be a simplicial complex on $[m]$. Suppose $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{a,b\}$. Then there exists an $i\in[m]$ such that
$\{a,b\}=\{i,i^{\prime}\}.$
Proof.
Denote by $\Gamma=\mathrm{sk}^{1}(\mathrm{Bier}(K))$. Since $a$ and $b$ are not linked by an edge in $\mathrm{Bier}(K)$, one and only one of the following three possibilities takes place:
  • $\{a,b\}\in\mathrm{MF}(K)$;
  • $\{a,b\}\in\mathrm{MF}(K^{\vee})$;
  • $\{a,b\}=\{i,i^{\prime}\}$ for some $i,1\leq i\leq m$.
However, in the first case, one has:
$\phi(a)\in\operatorname{Link}_{\Gamma}(b)\text{ and }\phi(a)\notin\operatorname{Link}_{\Gamma}(a)$
and similarly in the second case, one has:
$\phi^{-1}(a)\in\operatorname{Link}_{\Gamma}(b)\text{ and }\phi^{-1}(a)\notin\operatorname{Link}_{\Gamma}(a).$
This shows that only the last possibility takes place. ∎
Proposition 3.9.
Let $K$ be a simplicial complex on $[m]$ and $i\in[m]$. Then $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{i,i^{\prime}\}$ if and only if $K$ is a weak cone with apex $i$ and $K^{\vee}$ is a weak cone with apex $i^{\prime}$.
Proof.
If $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{i,i^{\prime}\}$, then $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, since they are subcomplexes. On the other hand, if $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, for all non-ghost vertices $j$ and $j^{\prime}$, $j\neq i$, all simplices $\{i,j\}$ and $\{i^{\prime},j^{\prime}\}$ are in $\mathrm{Bier}(K)$ and, by the construction of the Bier sphere also all simplices $\{i,j^{\prime}\}$ and $\{i^{\prime},j\}$. ∎
Lemma 3.10.
Let $K$ be a simplicial complex on $[m]$ and $i\in[m]$. Then the following statements are equivalent.
  • (a)
    The complex $K$ is a cone with an apex $i$.
  • (b)
    The complex $K^{\vee}$ is a cone with an apex $i^{\prime}$.
  • (c)
    The Bier sphere $\mathrm{Bier}(K)$ is a suspension with a vertex pair $\{i,i^{\prime}\}$.
Proof.
The equivalence (a) $\Leftrightarrow$ (b) follows from the observation that a vertex $i$ in $K$ belongs to each element in $\mathrm{M}(K)$ if and only if $i^{\prime}$ does not belong to any element in $\mathrm{MF}(K^{\vee})$ if and only if $i^{\prime}$ belongs to each element in $\mathrm{M}(K^{\vee})$.
The implication (a) $\Rightarrow$ (c) follows from [22, Lemma 3.3 (c)], where it was shown that
$\mathrm{Bier}(\operatorname{Cone}_{i}(K))=\Sigma_{\{i,i^{\prime}\}}(\mathrm{Bier}(K)).$
The implication (c) $\Rightarrow$ (a) holds, since each maximal simplex in $\mathrm{Bier}(K)$ belonging to $2^{[m]}$ contains the vertex $i$, hence $K=\mathrm{Bier}(K)\cap 2^{[m]}$ is a cone with apex $i$, see Figure 5. ∎
The example in the following figure is the motivation for the next lemma.
222233331111
[…]
2222333311114444
[…]
Figure 5. On the left is the graph $\Gamma_{6}$, which is self-dual, and its Bier sphere, which is the $6$-cycle. On the right is a cone over $\Gamma_{6}$, which is also self-dual, and its Bier sphere, which is a suspension over $\mathrm{Bier}(\Gamma_{6})$.
Lemma 3.11.
Let $K$ be a simplicial complex on $[m]$, $m\in V(K)$ and $L=\operatorname{Del}_{K}(m)$. Then one has:
$K=\operatorname{Cone}_{m}(L)\Longleftrightarrow\mathrm{Bier}(L)=\mathrm{Bier}(K)\cap 2^{[m-1]\sqcup[(m-1)^{\prime}]}.$
Proof.
The implication $\Rightarrow$ follows directly from [22, Lemma 3.3 (c)]:
$K=\operatorname{Cone}_{m}(L)\Rightarrow\mathrm{Bier}(K)=\Sigma_{\{m,m^{\prime}\}}(\mathrm{Bier}(L))\Rightarrow\mathrm{Bier}(L)=\mathrm{Bier}(K)\cap 2^{[m-1]\sqcup[(m-1)^{\prime}]}.$
Let us prove the implication $\Leftarrow$. Suppose $I\in\mathrm{M}(K)$ and $m\notin I$. Then $I\in\mathrm{M}(L)$, since $L$ is the deletion of $m$ from $K$. Therefore $m^{\prime}\in[m^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(K))$ and hence
$[(m-1)^{\prime}]\setminus I^{\prime}\notin\mathrm{MF}(\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}),$
by the definition of a minimal non-face. On the other hand, by Alexander duality, we have:
$[(m-1)^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(L)).$
Thus, $\mathrm{Bier}(L)$ and $\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}$ are different simplicial complexes on $[m-1]\sqcup[(m-1)^{\prime}]$, which contradicts our assumptions. This finishes the proof. ∎
Lemma 3.12.
Let $K$ be a simplicial complex on $[m]$. If either $f_{0}(K)$ or $f_{0}(K^{\vee})$ is less than $m$, then
$\chi(\mathrm{Bier}(K))=m-1$
if and only if one of the following combinatorial equivalences holds:
  • $K=\Delta_{[m-1]}$;
  • $K=\Delta_{[m-2]}$;
  • $K=\Delta_{[m-1]}\cup\Delta_{[m-2]\cup\{m\}}$.
In all those cases, one has:
$\mathrm{Bier}(K)=\mathrm{Bier}(\Delta_{[m-3]}\ast\Gamma_{4})=\Sigma^{m-3}(Z_{4}).$
Proof.
Suppose $\chi(\mathrm{Bier}(K))=m-1$ and let $c\colon V(\mathrm{Bier}(K))\to C=\{c_{1},c_{2},\ldots,c_{m-1}\}$ be a proper coloring. Without loss of generality, we can assume $f_{0}(K)<m$ and $\{m\}\notin K$. Therefore $[(m-1)^{\prime}]\in K^{\vee}$ and $C=\{c(1^{\prime}),\ldots,c((m-1)^{\prime})\}$. By Proposition 3.1 it follows that $\chi(K^{\vee})=m-1$.
If $\{m^{\prime}\}\in K^{\vee}$, then $m^{\prime}$ can not be linked by an edge to each of the vertices $1^{\prime},2^{\prime},\ldots,(m-1)^{\prime}$ and moreover:
$c(i)=c(i^{\prime})\text{ for all }1\leq i\leq f_{0}(K),$
since $\{i,j^{\prime}\}\in\mathrm{Bier}(K)$ for each $1\leq i\leq f_{0}(K)$ and each $1^{\prime}\leq j^{\prime}\leq m^{\prime}$ with $i\neq j$. For $j=m$ this implies that
$c(m^{\prime})\notin\{c(1^{\prime}),\ldots,c((f_{0}(K))^{\prime})\}.$
Thus, we conclude that $f_{0}(K)\leq m-2$, and there exists $j\in\{f_{0}(K)+1,\ldots,m-1\}$ such that $\{j^{\prime},m^{\prime}\}\not\in K^{\vee}$, since $\chi(K^{\vee})=m-1$. Furthermore, the vertex $j^{\prime}$ must be unique in this case, since having more than one such $j^{\prime}$ would imply the existence of two different $(m-3)$-dimensional simplices in $K$, which contradicts the condition $f_{0}(K)\leq m-2$.
Thus, we can assume that $\{(m-1)^{\prime},m^{\prime}\}\notin K^{\vee}$ and $\{j^{\prime},m^{\prime}\}\in K^{\vee}$ for all $21\leq j\leq m-$ 2. Hence finally $f_{0}(K)=m-2$ and moreover $K=\Delta_{[m-2]}$, with two ghost vertices $m-1$ and $m$. To complete the proper $(m-1)$-coloring $C$ of $\mathrm{Bier}(K)$ we set
$c(m^{\prime})=c((m-1)^{\prime}).$
If $\{m^{\prime}\}\notin K^{\vee}$, then $K^{\vee}=\Delta_{[(m-1)^{\prime}]}$ and hence $K=\Delta_{[m-1]}$. In this case, obviously, $\chi(\mathrm{Bier}(K))=m-1$ with coloring $C$ given by:
$i,i^{\prime}\mapsto c_{i}\text{ for all }1\leq i\leq m-1.$
This final statement follows from the definition of a Bier sphere, completing the proof of the first part. The second part follows from Example 2.11. ∎
Lemma 3.13.
Let $K$ be a simplicial complex on $[m]$ with $m\geq 4$ and $f_{0}(K)=f_{0}(K^{\vee})=m$. Then we have:
$\chi(\mathrm{Bier}(K))=m-1\Rightarrow\mathrm{Bier}(K)\text{ is a weak suspension}.$
Proof.
Suppose $\mathrm{Bier}(K)$ is not a weak suspension and $\chi(\mathrm{Bier}(K))=m-1$. Let $c\colon V(\mathrm{Bier}(K))=[m]\sqcup[m^{\prime}]\to C$ be a proper coloring of $\mathrm{Bier}(K)$, where $|C|=m-1$.
By our assumption and Proposition 3.9, either $K$ or $K^{\vee}$ is not a weak cone. We can assume that $K$ is not a weak cone, and we may also assume that $\{1,2\}\notin K$. Then $\{3^{\prime},4^{\prime},\ldots,m^{\prime}\}\in K^{\vee}$ and its vertices have pairwise different colors in $C$; let us define $C^{\prime}=\{c(3^{\prime}),\ldots,c(m^{\prime})\}$.
Since both $\{1,j^{\prime}\},\{2,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{3,\ldots,m\}$ it implies that $c(1),c(2)\notin C^{\prime}$ and therefore $c(1)=c(2)=:\alpha$, since $|C^{\prime}|=m-2$ and $|C|=m-1$. Thus, we have $C=C^{\prime}\sqcup\{\alpha\}$.
Suppose either $1$ or $2$ is not an isolated vertex of $K$. Without loss of generality, we can assume that $\{1,3\}\in K$. It implies that $c(3)\in C^{\prime}$. As $\{3,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{4,\ldots,m\}$ we get that $c(3)=c(3^{\prime})$.
Observe that $\{1,2^{\prime}\},\{1^{\prime},2\}\in\mathrm{Bier}(K)$ and hence $c(1^{\prime}),c(2^{\prime})\neq\alpha$. It follows that there are vertices $i^{\prime},j^{\prime}\in\{3^{\prime},4^{\prime},\ldots,m^{\prime}\}\in K^{\vee}$ such that $c(1^{\prime})=c(i^{\prime})$ and $c(2^{\prime})=c(j^{\prime})$. Hence the corresponding 2-element sets are non-edges in $K^{\vee}$ and Alexander duality yields:
$[m]\setminus\{1,i\},[m]\setminus\{2,j\}\in K.$
If $i\neq j$ the latter inclusions imply that $\{2,j\},\{1,i\}\in K$ and so $c(i),c(j)\neq\alpha$. Therefore
$c(i)=c(i^{\prime})=c(1^{\prime})\text{ and }c(j)=c(j^{\prime})=c(2^{\prime}),$
a contradiction, since $\{1^{\prime},i\},\{2^{\prime},j\}\in\mathrm{Bier}(K)$.
Thus, $i=j$ and $c(1^{\prime})=c(i^{\prime})=c(j^{\prime})=c(2^{\prime})=:\beta$ and
$[m]\setminus\{1,i\},[m]\setminus\{2,i\},\text{ and }[m]\setminus\{1,2\}\in K,$
since $\{1^{\prime},i^{\prime}\},\{2^{\prime},i^{\prime}\},\text{ and }\{1^{\prime},2^{\prime}\}\notin K^{\vee}$.
If $c(i)\neq\alpha$, then $c(i)\in C^{\prime}$ and so $c(i)=c(i^{\prime})$ by definition of Bier sphere. Then $\{1^{\prime},i\}\in\mathrm{Bier}(K)$ and $c(1^{\prime})=c(i^{\prime})=c(i)$, see above, a contradiction. Thus, $c(i)=\alpha$.
This implies that $\{1,i\},\{2,i\}\notin K$ and given $x$ such that $3\leq x\neq i\leq m$ we have
$x\in[m]\setminus\{1,i\},[m]\setminus\{2,i\},[m]\setminus\{1,2\}.$
It follows that $c(x)=c(x^{\prime})$ for all $3\leq x\leq m,x\neq i$.
Finally, our coloring goes as follows:
$c(1)$ $=c(2)=c(i)=\alpha\not\in C^{\prime},$
$c(1^{\prime})$ $=c(2^{\prime})=c(i^{\prime})=\beta\not\in C^{\prime},$
$c(x)$ $=c(x^{\prime})\text{ for }3\leq x\neq i\leq m.$
However, in such a Bier sphere, the pairs $\{j,j^{\prime}\}$ for each $j\in[m]\setminus\{1,2,i\}$ can be taken as a weak suspension vertex pair and therefore $\mathrm{Bier}(K)$ is a weak suspension, which contradicts the assumptions of the statement.
Now, suppose both $1$ and $2$ are isolated vertices of $K$. By Alexander duality, it follows that
$[m^{\prime}]\setminus\{1^{\prime},i^{\prime}\},[m^{\prime}]\setminus\{2^{\prime},i^{\prime}\}\in K^{\vee}\text{ for each }i\in\{3,\ldots,m\}.$
Recall that we also have $[m^{\prime}]\setminus\{1^{\prime},2^{\prime}\}\in K^{\vee}$. Therefore, since $m\geq 4$ by the assumptions of the statement, we have $c(1^{\prime}),c(2^{\prime})\notin C^{\prime}=C\setminus\{\alpha\}$, which implies that $c(1^{\prime})=\alpha=c(2^{\prime})$, a contradiction as $\{1^{\prime},2\},\{1,2^{\prime}\}\in\mathrm{Bier}(K)$ and $\alpha=c(1)=c(2)$. Therefore, $\mathrm{Bier}(K)$ is a weak suspension, which finishes the proof. ∎
Lemma 3.14.
Let $K$ be a simplicial complex on $[m]$. Suppose $\mathrm{Bier}(K)$ is a weak suspension, which is not a suspension. Then we have:
$\chi(\mathrm{Bier}(K))=m.$
Proof.
Due to Lemma 3.8, we can assume that $\{1,1^{\prime}\}$ is a weak suspension vertex pair for $\mathrm{Bier}(K)$.
Since $\mathrm{Bier}(K)$ is not a suspension, there exists a simplex $\sigma\in\operatorname{Del}_{K}(1)$ such that $\{1\}\cup\sigma\notin K$. Then, by definition of Alexander duality, its complement $[2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime}\in K^{\vee}$ and therefore, by definition of Bier sphere, $\tau:=([2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime})\cup\sigma$ is a maximal simplex in $\mathrm{Bier}(K)$.
By Proposition 3.9, $K$ is a weak cone with apex $1$, hence the vertex $1$ is linked by an edge in $\mathrm{Bier}(K)$ to each vertex of $\sigma\in K$ and therefore, to each vertex of $\tau$. Since $|\tau|=m-1$, it follows that the 1-skeleton of the full subcomplex of $\mathrm{Bier}(K)$ on the vertex set $\{1\}\cup\tau$ is a complete graph with $m$ vertices. This yields $\chi(\mathrm{Bier}(K))=m$, by Proposition 3.1, which finishes the proof. ∎
The next result characterizes the case when $\chi(\mathrm{Bier}(K))=m-1$ for a simplicial complex $K$ on $[m]$.
Lemma 3.15.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ and $m\geq 3$.
Then $\chi(\mathrm{Bier}(K))=m-1$ if and only if $K$ is obtained from either $G_{4}$, or $\Gamma_{4}$, or $\Gamma_{6}$ in Figure 2 by iterative application of the following two simplicial operations:
  • taking a cone;
  • taking Alexander dual.
Proof.
To prove the implication $\Leftarrow$, we observe that by [22, Lemma 3.3 (c)] one has:
$\mathrm{Bier}(\Delta\ast\Gamma_{4})=\mathrm{Bier}(\Delta\ast G_{4})=\Sigma^{m-3}(Z_{4})\text{ and }\mathrm{Bier}(\Delta\ast\Gamma_{6})=\Sigma^{m-3}(Z_{6}).$
Coloring each time the vertex pair of the suspension into a new color, we get
$\gamma(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{4})\text{ or }\gamma(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{6}),$
respectively. Both right hand sides are equal to $m-1$. Application of Proposition 3.1 finishes the proof of this implication.
To prove the implication $\Rightarrow$, we argue by induction on $m$. For $m\leq 4$ the statement is true by the complete description of combinatorial Bier spheres in dimensions 1 and 2, see [21, Theorem 2.16]. Suppose the statement holds for a fixed $m\geq 4$, and let us show it is true when $K$ is a simplicial complex on $[m+1]$. Due to Lemma 3.12 it remains to consider the case when both $K$ and $K^{\vee}$ have no ghost vertices:
$f_{0}(K)=f_{0}(K^{\vee})=m+1.$
By our assumption, $\chi(\mathrm{Bier}(K))=m$ and therefore, by Lemma 3.13, $\mathrm{Bier}(K)$ is a weak suspension. Then, by Lemma 3.14, $\mathrm{Bier}(K)$ is a suspension.
Due to Lemma 3.8 we can assume that $\{m+1,(m+1)^{\prime}\}$ is a vertex pair of $\mathrm{Bier}(K)$, hence, by Lemma 3.10, $m+1$ and $(m+1)^{\prime}$ are the apexes of $K$ and $K^{\vee}$, respectively. Consider $R:=\operatorname{Del}_{K}(m+1)$ as a simplicial complex on $[m]$. Then $K=\operatorname{Cone}_{m+1}(R)$ and due to Lemma 3.11 one has:
$\mathrm{Bier}(R)=\mathrm{Bier}(K)\cap 2^{[m]\cup[m^{\prime}]}.$
Since $\chi(\mathrm{Bier}(K))=m$ and $\mathrm{Bier}(K)$ is a suspension over $\mathrm{Bier}(R)$, by Lemma 3.7 we get:
$\chi(\mathrm{Bier}(R))=\chi(\mathrm{Bier}(K))-1=m-1.$
By the inductive assumption applied to $\mathrm{Bier}(R)$, we obtain that $R$ is obtained from either $G_{4}$, or $\Gamma_{4}$, or $\Gamma_{6}$ by iterative application of the following two simplicial operations: taking a cone and taking Alexander dual.
It remains to recall that $K=\operatorname{Cone}_{m+1}(R)$, which finishes the proof. ∎
In what follows, we use the notation $P_{K}$ for the combinatorial simple polytope such that
$\mathrm{Bier}(K)=\partial P^{*}_{K},$
when $\mathrm{Bier}(K)$ is a polytopal sphere. Now we are ready to introduce a complete description of chromatic numbers for Bier spheres, which summarizes the results of this section.
Theorem 3.16.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then
  1. (a)
    $m-1\leq\chi(\mathrm{Bier}(K))\leq m$;
  2. (b)
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next two cases holds:
    • $P_{K}=I^{m-1}$;
    • $P_{K}=I^{m-3}\times P_{6}$.
Proof.
The case $m=2$ is clear. In case $m\geq 3$, statement (a) follows directly from Proposition 3.1, while statement (b) is a direct consequence of Lemma 3.15. ∎
4. Buchstaber numbers and chordality of Bier spheres
In this section, we discuss Buchstaber numbers of Bier spheres and their chordality. These combinatorial invariants and combinatorial property are closely related to certain important topological properties of polyhedral products studied in the framework of toric topology [9, 10].
Definition 4.1.
Let $K$ be a simplicial complex on $[m]$ with $m$ geometrical vertices. We call the complex Buchstaber number of $K$ the maximal integer $r$ such that there exists a characteristic map
$\Lambda\colon[m]\to\mathbb{Z}^{m-r}$
and we denote it by $s(K)$; that is, each simplex of $K$ is mapped to a part of a lattice basis of $\mathbb{Z}^{m-r}$.
Similarly, given a prime number $p\in\mathbb{N}$, we call the mod $p$ Buchstaber number of $K$ the maximal integer $r$ such that there exists a mod $p$ characteristic map
$\Lambda_{p}\colon[m]\to\mathbb{Z}_{p}^{m-r}$
and we denote it by $s_{p}(K)$; that is, each simplex of $K$ is mapped to a linearly independent set of vectors in the $\mathbb{Z}_{p}$-vector space $\mathbb{Z}_{p}^{m-r}$.
Remark.
Note that this definition is equivalent to the one given in [6, Definition 5.3]; moreover, in the case $p=2$, it is equivalent to the one of $s_{\mathbb{R}}(K)$ as the maximal rank of a real toric subgroup in $\mathbb{Z}_{2}^{m}$ acting freely on the real moment-angle-complex $\mathcal{R}_{K}$ and similarly in the complex case, see [6, Definition 5.1, Remark 5.5].
It follows from the above definition that both $s(K)$ and $s_{p}(K)$, for any prime $p\in\mathbb{N}$, are combinatorial invariants of $K$. The real Buchstaber number was introduced and studied in [14], where it was proven that
$1\leq s(K)\leq s_{\mathbb{R}}(K)\leq m-n,$
for any simplicial complex $K$ with $f_{0}(K)=m$ and $\dim(K)=n-1$, see also [1, Proposition 3] for an alternative proof. The mod $p$ Buchstaber invariant was introduced and studied in [6], where it was shown, see [6, Corollary 5.8], that
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n,\text{ for any prime $p\in\mathbb{N}$}.$
Since $\dim(\mathrm{Bier}(K))=m-2$, applying these inequalities and Theorem 3.16 to a Bier sphere, one has:
$f_{0}(\mathrm{Bier}(K))-m\leq s(\mathrm{Bier}(K))\leq s_{p}(\mathrm{Bier}(K))\leq f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
Our next result shows that the last two inequalities above are, in fact, equalities for any Bier sphere.
In [21, Theorem 3.4], the real and complex Buchstaber invariants of $\mathrm{Bier}(K)$ were computed assuming that $K$ is considered alongside with its ghost vertices. Now, we are going to state and prove the general formulae for all Buchstaber numbers of $\mathrm{Bier}(K)$, when only geometrical vertices of $K$ are taken into account.
Theorem 4.2.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then we have
$s(\mathrm{Bier}(K))=s_{p}(\mathrm{Bier}(K))=f_{0}(K)-f_{m-2}(K)+1,\text{ for any prime }p\in\mathbb{N}.$
Proof.
First, let us compute the number of geometrical vertices of $\mathrm{Bier}(K)$ in terms of the $f$-vector of $K$. By [21, Corollary 2.7], the Bier sphere $\mathrm{Bier}(K)$ has $(m-f_{0}(K))+f_{m-2}(K)$ ghost vertices, which is just the sum of the numbers of ghost vertices in $K$ and in its Alexander dual complex $K^{\vee}$. Hence the number of geometrical vertices of the Bier sphere equals
$|V(\mathrm{Bier}(K))|=f_{0}(\mathrm{Bier}(K))=2m-(m-f_{0}(K)+f_{m-2}(K))=m+f_{0}(K)-f_{m-2}(K).$
Denote by $\{e_{1},e_{2},\ldots,e_{m-1}\}$ the standard basis of $\mathbb{Z}^{m-1}$ and consider the map
$\phi\colon[m]\sqcup[m^{\prime}]\to\mathbb{Z}^{m-1}\text{ such that }\phi\colon i,i^{\prime}\mapsto e_{i},\text{ for all }1\leq i\leq m-1,\text{ and }\phi\colon m,m^{\prime}\mapsto e_{1}+e_{2}+\ldots+e_{m-1}.$
Let us define $\Lambda_{(p)}$ to be the restriction of $\phi$ to $V(\mathrm{Bier}(K))$. Then $\Lambda_{(p)}$ is a characteristic map for $\mathrm{Bier}(K)$ both in the complex and mod $p$ case, for each prime $p\in\mathbb{N}$. This follows from the fact that maximal simplices of $\mathrm{Bier}(K)$ have cardinality $(m-1)$ and do not contain a pair $\{i,i^{\prime}\}$, for any $1\leq i\leq m$.
Set $r=f_{0}(\mathrm{Bier}(K))-(m-1)$. By the definition of a Buchstaber number, this shows that
$s_{p}(K)\geq s(K)\geq r=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
On the other hand, we always have:
$s(K)\leq s_{p}(K)\leq f_{0}(\mathrm{Bier}(K))-(\dim(\mathrm{Bier}(K))+1)=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
Thus one gets the following equality
$s(K)=s_{p}(K)=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
Finally, the right hand side equals $(m+f_{0}(K)-f_{m-2}(K))-(m-1)=f_{0}(K)-f_{m-2}(K)+1$. Q.E.D. ∎
Now we turn to the discussion of chordality for Bier spheres.
Definition 4.3.
We say that a graph $\Gamma$ is chordal if there are no induced cycles of length greater than $3$ in $\Gamma$; in other words, if $\Gamma_{I}=Z_{k}$, then $k=3$; such a full subcomplex is called chordless. Finally, we call a simplicial complex $K$ chordal if $\mathrm{sk}^{1}(K)$ is a chordal graph.
Definition 4.4.
We write $P=\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 2,k\geq 0$ and say that $P$ is a truncation polytope if $P$ is combinatorially equivalent to a convex simple $n$-dimensional polytope obtained from $\Delta^{n}$ by a sequence of $k$ vertex truncations by means of hyperplanes in general position.
Figure 6. Some examples of types $\mathrm{vc}^{0}(\Delta^{3})$, $\mathrm{vc}^{1}(\Delta^{3})$, $\mathrm{vc}^{2}(\Delta^{3})$, $\mathrm{vc}^{3}(\Delta^{3})$, and $\mathrm{vc}^{4}(\Delta^{3})$.
Remark.
Note that the polytope $Q$ combinatorially dual to a truncation polytope $P$ is combinatorially equivalent to a convex simplicial polytope obtained from $\Delta^{n}$ by a sequence of $k$ stellar subdivisions in facets. Such a polytope is called a stacked polytope. Observe that the combinatorial types of $P$ and $Q$ depend on the vertices truncated and facets subdivided, respectively, when $n\geq 3$ and $k\geq 3$. Namely, there is only one combinatorial type of $\mathrm{vc}^{0}(\Delta^{3})$, $\mathrm{vc}^{1}(\Delta^{3})$, and $\mathrm{vc}^{2}(\Delta^{3})$, whereas there are 3 combinatorially different realizations of $\mathrm{vc}^{3}(\Delta^{3})$ (see Figure 7) and 7 combinatorially different realizations of $\mathrm{vc}^{4}(\Delta^{3})$ [16, Table 8].
Figure 7. Examples of types of $\mathrm{vc}^{3}(\Delta^{3})$. The last two polytopes are combinatorially equivalent.
Proposition 4.5.
Let $P=\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 3,k\geq 0$ be a truncation polytope. Then $K_{P}$ is chordal.
Proof.
Since for $P=\Delta^{n},n\geq 3$ the statement is true, we can fix $n$ and use induction on $k\geq 0$.
To make the induction step, suppose $P=\mathrm{vc}^{k+1}(\Delta^{n})$ and $n\geq 3,k\geq 0$. Then the dual stacked polytope $Q$ was obtained by performing a stellar subdivision with apex $v$ in a facet $\sigma:=\{i_{1},\ldots,i_{n}\}$ of a stacked polytope $Q^{\prime}$ which is dual to a certain simple polytope of the type $\mathrm{vc}^{k}(\Delta^{n})$.
Suppose $S$ was an induced $p$-cycle in the graph $\mathrm{sk}^{1}(\partial Q)$ and $p\geq 4$. We have two possible cases.
If $v\notin S$, then $S$ was an induced cycle in $\partial Q^{\prime}$, which is chordal by inductive assumption, and hence $p=3$, a contradiction. Indeed, since $\sigma$ was a simplex in $\partial Q^{\prime}$, the induced chordless cycle $S$ of length $p\geq 4$ could intersect it in no more than two vertices. Then $S$ remains a full subcomplex in $\partial Q$ after the stellar subdivision in $\sigma$ was performed.
If $v\in S$, then at least two vertices, $u$ and $w$, of $\sigma$ must belong to the cycle $S$. Therefore, $S$ must contain all the three edges, $\{u,v\},\{u,w\}$, and $\{v,w\}$. This shows that $S$ is not a chordless cycle, a contradiction. This finishes the proof. ∎
We are ready to state the result, which provides the classification of all chordal Bier spheres in dimensions greater than one.
Lemma 4.6.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 4$. Then $\mathrm{Bier}(K)$ is chordal if and only if either $K$ or $K^{\vee}$ contains no edges.
Proof.
The implication $\Leftarrow$ follows immediately from Proposition 4.5. Indeed, the case of a void complex is obvious. Suppose $k>0$ and $K^{\vee}$ is generated by its vertex set $\{1^{\prime},2^{\prime},\ldots,k^{\prime}\}$. By definition of a Bier sphere, it implies that $\mathrm{Bier}(K)$ is obtained from $\partial\Delta_{[m]}$ by stellar subdivisions of the facets $[m]\setminus\{i\}$ by vertices $i^{\prime}$ for $1\leq i\leq k$, see Figure 6. We obtain a polytopal Bier sphere, whose Bier polytope $P_{K}$ is of the type $\mathrm{vc}^{k}(\Delta^{m-1})$.
To prove the implication $\Rightarrow$, suppose that $\mathrm{Bier}(K)$ is chordal and $\mathrm{sk}^{1}(K)$ contains an edge. Without loss of generality, we can assume that $\{1,2\}\in K$.
Then $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$, otherwise $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. It follows that if $\{i\}\notin K$ with $i\geq 3$, then Alexander duality would imply that $\{1^{\prime},2^{\prime}\}\subset[m^{\prime}]\setminus\{i^{\prime}\}\in K^{\vee}$, which contradicts $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$. This yields that $f_{0}(K)=m$ and $K$ has no ghost vertices.
Now, we are going to show that both $\{1,i\}$ and $\{2,i\}$ are in $K$, for each $i$ with $3\leq i\leq m$. Indeed, otherwise one of the following three cases occurs. If there exists an $i,3\leq i\leq m$ such that $\{1,i\},\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 5-cycle on the vertices $\{1,2,1^{\prime},i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. If there exists an $i,3\leq i\leq m$ such that $\{1,i\}\notin K$ and $\{2,i\}\in K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. Finally, if there exists an $i,3\leq i\leq m$ such that $\{1,i\}\in K$ and $\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},i\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
Next, observe that $\{i,j\}\in K$, for every pair of vertices $i,j$ such that $3\leq i\neq j\leq m$. Indeed, if $\{i,j\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,i,1^{\prime},j\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
Finally, it follows that $K^{\vee}$ contains no edges. Indeed, as was already shown above, $\mathrm{sk}^{1}(K)$ is a complete graph with $m$ vertices. Suppose there was an edge $\{i^{\prime},j^{\prime}\}\in K^{\vee}$ with $1\leq i\neq j\leq m$. Then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{i,j,i^{\prime},j^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. This finishes the proof. ∎
We summarize our results on chordality for Bier spheres in the next statement.
Theorem 4.7.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then $\mathrm{Bier}(K)$ is chordal if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next three cases holds:
  • $P_{K}=\Delta^{1}$, where $m=2$;
  • $P_{K}=\Delta^{2}$, where $m=3$;
  • $P_{K}=\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
Proof.
In case $m=2$ the unique 00-dimensional Bier sphere is a chordal complex. In case $m=3$ the only chordal $1$-dimensional Bier sphere is the boundary of a triangle, see Figure 1. If $m\geq 4$, then the application of Lemma 4.6 finishes the proof. ∎
Acknowledgements
The authors are grateful to Djordje Baralić, Matvey Sergeev, and Rade Živaljević for numerous fruitful discussions, valuable comments and suggestions.
Limonchenko was supported by the Serbian Ministry of Science, Innovations and Technological Development through the Mathematical Institute of the Serbian Academy of Sciences and Arts.
Vavpetič was supported by the Slovenian Research and Innovation Agency program P1-0292 and the grant J1-4031.